36 microscopelog
Ian C. Clark edited this page 4 years ago

21/02/2019: @icc20 SLM775 before:

Var 2D 3D
OffX 11 29
OffY 22 74
GrX 4.64/mm 4.77/mm
GrY 1.35/mm 1.54/mm
Rad 5 6.20
Phase -2 1
Rot 0 90

SLM775 after:

Var 2D 3D
OffX 7 29
OffY 30 74
GrX 4.80/mm 4.77/mm
GrY 1.56/mm 1.54/mm
Rad 5 6.20
Phase -2 1
Rot 0 90

20/11/2018: @icc20

New software. The default template for the new software does not work. If one loads a template (such as the confocal template) things start to work more as expected. Somewhere we have documentation on how to set the default for all users, I'll just have to search it out. Perhaps:

click on “measurement” (toolbar) -> “make settings default”.

Obvious changes are the new easyCommander UI and a more adaptable (in some ways clearer) pixel steps UI. The default operating power for the STED lasers also seems to have changed, it's much lower.

28/08/2018: @icc20 @mol21 Tested replacement shutter. New shutter works. Noted poorly terminated shutter cable.

Abberior informed, and asked for electronic copy of documentation.

17/08/2018: @icc20 Turned system on ready for support call with Abberior to try to solve shutter status problem.

Estabilished: Shutter interlock display on equipment box agrees with software -> not an equipment/pc interface problem No obvious loose cables outside the equipment box.

Actions: @icc20 will attempt to fully trace shutter cables. Abberior will sent replacement shutter to aid in diagnosis Abberior will investigate whether this can be a user replaceable part

08/08/2018: ICC Checked after changing compressor. 2D y-grating to 1.40/mm x-grating 4.79/mm

07/08/2018: ICC Replaced borrowed compressor.

06/08/2018: @mol21 775nm SLM needed realignment changed 2D STED y-grating to: 1.12/mm

03/08/2018: ICC

Logged in with unprivileged network account. System started, but safety shutter remained closed. Switching user to local unprivileged account, leaving system turned on. There is a live dialog "system_status" indicating the safety shutter status.

Suspect there is a logic in there somewhere. Once the settings in the measurement are correct and image can be acquired, going to "Measurement > Make Settings Default" saves the info.

In "SCAN" mode, changing the FL shutter and rotating through the lightpath (top right button on touch-screen) seems to work reliably, even when using a network user.

01/08/2018: MOL

problems with startup, frame was not recognised, had to re-initialise error, could not start measurement on ix83. switching between frame operation using panel and impector does not work smoothly still a problem, switching back and forth, starting new measurement, open/closing fluorescence shutter using panel in random order allows to overcome this error

Error Screenshot 1

Error Screenshot 2

changes of alignment of 775SLM grating, 2D y is now 1.12 3D: 4.72 and 1.34

aligned 561nm laser to the other excitation lasers using beads, was off to the bottom right

26/7/2018: MOL

Login as user, loaded template for red STED (640 STED and confocal), STED worked but misaligned changed 775 SLM grating 2D: x: 4.79 y: 1.31 3D: x: 5.07 y: 1.34

595 SLM grating: 2D x: 5.09 y: 1.19 3D x: 4.99 y: 1.42

24/7/2018: MOL

Started up the microscope without any problem, using IX83, Used Tetraspec beads and could see immediate STED with STAR 635 channel, STED laser at 15%, exc and STED slightly misaligned, but overlap was still in the PSF, changed grating y for 775nm STED for 2D (from 1.15/mm to 1.23/mm) for 3D (from 1.14/mm to 1.28/mm)